OpenStudy (anonymous):

The simple formula for the nth term of an arithmetic sequence is an = 4n + 16. What is the explicit formula corresponding to the simple formula? A.an = 4 + (n - 1)(16) B.an = 16 + (n - 1)(4) C.an = 16 + 4n D.an = 20 + (n)(4) E.an = 20 + (n - 1)(4)

OpenStudy (anonymous):

@Robert136

OpenStudy (welshfella):

an = a1 + (n - 1)d

OpenStudy (welshfella):

where d = common difference and a1 = first term

OpenStudy (anonymous):

so its a???

OpenStudy (welshfella):

in your case d = 4

OpenStudy (anonymous):

mhm

OpenStudy (welshfella):

no

OpenStudy (anonymous):

ik its not a and b

OpenStudy (welshfella):

the coefficient of n is d in the explicit formula so that makes d = 4 but what value will a1 have?

OpenStudy (welshfella):

why do you say its not b?

OpenStudy (welshfella):

oh i see - no its not b

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

im i right>??

OpenStudy (welshfella):

simplif each of the options and see if you get 4n + 16

OpenStudy (welshfella):

no

OpenStudy (anonymous):

how when its already 16+4n

OpenStudy (welshfella):

20 + (n - 1)4 = 20 + 4n - 4 = ?

OpenStudy (anonymous):

the -4 will cancel out

OpenStudy (anonymous):

so it will be left with 2o+44n then you divide both by 4

OpenStudy (welshfella):

you need the explicit formula where you can easily pick out the first term and the common difference

OpenStudy (welshfella):

20 - 4 = 16

OpenStudy (anonymous):

so its d

OpenStudy (welshfella):

theres only 1 left lol

OpenStudy (anonymous):

d

OpenStudy (welshfella):

e = 20 + (n - 1)4 = 10 + 4n - 4 = 4n + 16 so its e

OpenStudy (anonymous):

ok EXPLAIN PLZ

OpenStudy (welshfella):

* that is 20 + 4n - 4 in second line

OpenStudy (anonymous):

ik that

OpenStudy (welshfella):

- well it simplifies back to 4n + 16 and a1 + (n - 1)d is the explicit formula

OpenStudy (anonymous):

oooo lol thx

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

ill tag you one my other one